OpenStudy (anonymous):

Need some help, lots actually, with Trigonometric Identities. \[\frac{ \sec \Theta }{ \csc \Theta - \cot \Theta} - \frac{ \sec \Theta }{ \csc \Theta + \cot \Theta } = 2 \csc \Theta\] Have no idea where to start.

OpenStudy (welshfella):

is that right for the left side?

OpenStudy (anonymous):

Looks right, Here is what it is \[\frac{ \sec \Theta }{ \csc \Theta - \cot \Theta} - \frac{ \sec \Theta }{ \csc \Theta + \cot \Theta } = 2 \csc \Theta\]

OpenStudy (welshfella):

oh ok i made a mistake in first fraction

OpenStudy (welshfella):

hold on - let me rethink this one

OpenStudy (anonymous):

Sure, no problem. I have a list here that says \[\csc \Theta = \frac{ 1 }{ \sin \Theta }, \sec \Theta = \frac{ 1 }{ \cos \Theta }, \cot \Theta = \frac{ 1 }{ \tan \Theta }\] if that helps at all. Just not sure how to make them work.

OpenStudy (welshfella):

subtractingthe 2 fractions on LHS we get secO(cscO + cotO) - secO(cscO - cotO) --------------------------------- csc^2 O - cot^2 O the denominator = 1 so we have

OpenStudy (welshfella):

secO(cscO + cotO) - secO(cscO - cotO)

OpenStudy (welshfella):

simplifying this we get 2 sec O cot O

OpenStudy (anonymous):

I think I left out something in the original post. I am suppose to verify it.

OpenStudy (welshfella):

2 sec O cot O = 2 cos O 2 --- * ----- = --- = 2 csc O = RHS cos O sin O sin O

OpenStudy (welshfella):

- there yougo - this verifies it

OpenStudy (welshfella):

yes when verifying you need to use known identities to convert the more complex side of the identity to get the other side (LHS - left hand side, RHS = right hand side)

OpenStudy (welshfella):

did you follow that ok ?

OpenStudy (anonymous):

Not really, Trying to figure it out but i'm just not getting it

OpenStudy (welshfella):

there's a known identity csc^ x = 1 + cot^x when we rearrange this it becomes csc^2 x - cot^x = 1 i used that one to simplify the fraction

OpenStudy (welshfella):

I slso used sec x = 1/ cos x and cot x = cos x / sin x

OpenStudy (welshfella):

subtracting the fractions was just algebra

OpenStudy (anonymous):

I'm sorry just not understanding it for some reason. It's probably really simple once it's understood.

OpenStudy (welshfella):

yes do you see how i got csc^2 O - cot^2 O as the denominator?

OpenStudy (welshfella):

its simply the product of csc O - cot O and csc o + cot O

OpenStudy (welshfella):

like ( a + b)( a - b) = a^2 - b^2

OpenStudy (welshfella):

anyway sorry but i gotta go right now good luck